OpenStudy (anonymous):

how should I solve f(15) = 1/3 (15) - 2?

OpenStudy (anonymous):

you divide and subtract

OpenStudy (anonymous):

How do I solve it with the fraction though?

OpenStudy (anonymous):

fraction just means to divide it

OpenStudy (anonymous):

(1/3)15 = 5

OpenStudy (anonymous):

The answer would be 3 (: Plugged it in to my calculator.

OpenStudy (anonymous):

The fraction doesn't represent a division sign though... unless I'm missing something

jimthompson5910 (jim_thompson5910):

f(15) = 1/3 (15) - 2 f(15) = (1/3)* (15/1) - 2 f(15) = (1*15)/(3*1) - 2 f(15) = (15)/(3) - 2 f(15) = 5 - 2 f(15) = 3

OpenStudy (anonymous):

it is just defined that way. you'll have to take it as a division

OpenStudy (anonymous):

Okay. Even if it's 1 over 3? (Just to clarify.)

OpenStudy (anonymous):

if you go deep, 1/3 will be 3's inverse for product

jimthompson5910 (jim_thompson5910):

fractions are a division of two numbers ex: 2 divided by 3 = the fraction 2/3

OpenStudy (anonymous):

i don't know your knowledge background so i will stop. anyways 4*(1/2) means divide 4 by 2 15*(1/3) means divide 15 by 3

OpenStudy (anonymous):

this works only for the ones that look like 1/a

OpenStudy (anonymous):

I understand that, written on on my assessment it's written as a fraction. I understand that I wrote it as division, I didn't no any other way to do it.

OpenStudy (anonymous):

It's written as 1 -- 3